the average weight of randomly selected 28 adult males was 65kigs. with standard deviation of 2.2 kgs. find a 99% confidence interval


1
Expert's answer
2022-05-01T17:22:14-0400

The formula to calculate a confidence interval for a population mean is as follows:

"CI=\bar{x}\pm z\cdot\cfrac{s}{\sqrt{n}},"

where:

  • "\bar{x}=65" : sample mean
  • zthe chosen z-value, for a 99% confidence interval z = 2.576
  • "s=2.2" sample standard deviation
  • n = 28 sample size.

So,

"CI=65 \pm 2.576\cdot\cfrac{2.2}{\sqrt{28}}=\\=
65\pm1.07=(63.93, 66.07)."